OpenStudy (anonymous):

Solve each equation using factoring. Write the answer in set brackets in the blank like this with no spaces:{1,2}. x^2+7x+12=0

OpenStudy (anonymous):

what are 2 numbers whose product is 12 and whose sum is 7?

OpenStudy (anonymous):

10+2=12 , and 5+2=7 ??

OpenStudy (anonymous):

no the same 2 numbers?

OpenStudy (anonymous):

3 and 4 3+4=7 3*4=12 so x^2+7x+12=(x+4)(x+3) can you take it from here?

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

ok (x+4)(x+3)=0 => x+4=0 =>x=-4 or x+3=0 =>x=-3 {-4,-3}